OpenStudy (nuccioreggie):

Which of the following would be equivalent to 11 to the 8th power over 11 to the 3rd power? 11 to the 7th power over 11 to the 7th power 11 ⋅ 114 11 to the 0 power over 11 to the 5th power 110

OpenStudy (studygurl14):

\[\frac{ 11^8 }{ 11^3 }=11^5=161051\] None of the answer choices are correct...are you sure C is not 11th to the 0 power over 11 to the negative 5th power?

OpenStudy (nuccioreggie):

it is wat it say